# Pitch: AI-assisted Tagging of Deepfake Audio Calls using Challenge-Response

### v3

@@ 4.2. Challenges Dataset Curation
We carefully considered the ethical implications of creating and using deepfakes, including potential misuse and privacy concerns. To mitigate these risks, we implemented strict data handling protocols and enforcement to access the generated deepfakes.
@@
We developed the experiments using the Gorilla Experiment Builder (see Fig A.4).***We open our instrument at https://app.gorilla.sc/openmaterials/722500. The initial data collection involved 100 subjects, equally distributed between smartphone and desktop environments, yielding 18,600 original recordings. This dataset is comparable in scale to the well-known VCTK speech dataset (Yamagishi et al., 2019), which contains 110 speakers with 400 samples each (totaling 44,000 recordings), while offering significantly greater diversity. Our participant pool consisted of individuals aged 18-65, near equal gender distribution, and is multi-racial. This diverse demographic helped ensure a representative sample of speech patterns and characteristics.
@@
Three English-speaking evaluators—a native monolingual, a non-native multilingual, and a tie-breaker—verified each sample. This approach aimed to enhance inclusivity and mitigate demographic bias. We made compliance decisions using majority voting, resulting in an 84% mean compliance across 100 subjects and 89% across the 21 challenges. Fig. A.6 illustrates the subject-compliance distribution across the dataset and demographic distribution.
Choosing RTDF generators. Our analysis revealed that state-of-the-art Text-to-Speech (TTS) models such as TorToise (Betker, 2023) and Bark (suno-ai, 2023) excelled in content modulation but fell short in capturing speech nuances like emotions or prosody, significantly limiting their effectiveness against our challenges. Consequently, we shifted our focus to speech-to-speech systems, which convert source speech into the targeted speaker’s voice while preserving subtle speech elements.
Among the various voice converters available, we selected FREE-VC (Li et al., 2023b), PPG-VC (Liu et al., 2021), and StarGANv2-VC (Li et al., 2021) for our evaluation. We most thoroughly evaluate FREE-VC, as it is open-source and has any-to-any voice cloning capabilities, outperforming earlier models and matching recent advancements. A threat analysis study by Korshunov et al. (Korshunov et al., 2023) identified it as a high threat to speaker recognition systems, suggesting its potential as an imposter’s preferred choice. PPG-VC represented a prior state-of-the-art deepfake generator, while StarGANv2-VC was the subject of evaluation in previous work by Yasur et al. (Yasur et al., 2023).
@@
| Baseline | Val Acc. | Test Acc. | AUC | Precision | Recall | F1 | ||
| #1 | Static Mouth | 85.9 | 88.2 | 96.6 | 0.89 | 0.87 | 0.88 | |
| #2 | Cup mouth | 86.9 | 89.5 | 97.1 | 0.95 | 0.84 | 0.89 | |
| #3 | Whisper | 94.9 | 90.9 | 93.1 | 0.92 | 0.9 | 0.91 | |
| #4 | Hold nose | 43.4 | 90.0 | 92.8 | 0.94 | 0.85 | 0.90 | |
| #5 | High Pitch | 86.9 | 81.8 | 91.2 | 0.95 | 0.67 | 0.79 | |
| #6 | Low Pitch | 73.7 | 77.3 | 85.9 | 0.75 | 0.83 | 0.78 | |
| #7 | Sing | 87.9 | 90.0 | 96.8 | 0.89 | 0.91 | 0.9 | |
| #8 | Speak Loudly | 86.9 | 70.5 | 74.8 | 0.73 | 0.65 | 0.69 | |
| #9 | Speak softly | 85.9 | 79.1 | 84.7 | 0.76 | 0.85 | 0.8 | |
| #10 | Read quickly | 98.0 | 83.2 | 90.7 | 0.92 | 0.73 | 0.81 | |
| #11 | Read Slowly | 89.9 | 79.1 | 87.5 | 0.8 | 0.77 | 0.79 | |
| #13 | Accent | 72.7 | 79.5 | 85.9 | 0.81 | 0.77 | 0.79 | |
| #14 | Emotions | 89.9 | 88.6 | 95.6 | 0.98 | 0.79 | 0.87 | |
| #16 | Cough/Whistle | 90.9 | 96.7 | 99.3 | 0.82 | 0.82 | 0.82 | |
| #17 | Clap | 84.7 | 94.1 | 98.8 | 0.95 | 0.93 | 0.94 | |
| #18 | Cross-talk | 91.5 | 98.8 | 100.0 | 1.0 | 0.96 | 0.98 | |
| #19 | Instr. Playback | 95.7 | 98.1 | 99.6 | 0.98 | 0.96 | 0.97 | |
| #20 | Lyric Playback | 100.0 | 98.1 | 100.0 | 0.94 | 1.0 | 0.97 | |
| Average (across all) | 86.0 | 87.5 | 93.1 | 0.89 | 0.82 | 0.85 | ||
| Heuristic-based Compliance Check | ||||||||
| #15 | Question | 92.9 | Transcribed sentence has a question mark | |||||
@@ 5. Evaluation
5.1. RQ1. Machine Evaluation
We evaluate the effectiveness of our framework against voice-cloning attacks using an automated authentication system. An impersonator’s goal is to mimic a target individual during a call by:
Adhering to the challenge in terms of task and content,
@@
Conversely, our challenges aim to disrupt these goals. To assess challenge performance, we quantify the disruption caused. However, incorporating speaker identity verification would require a reference sample, making the evaluation intrusive (needing a reference sample of the target’s voice) and thus, impractical. Therefore, we focus on three non-intrusive criteria:
These criteria were selected for their ability to comprehensively assess voice sample quality and authenticity without reference samples, maintaining the non-intrusive nature of our evaluation.
@@
We fine-tuned each Wav2Vec2 model for 30 epochs with a batch size of 64 and a learning rate of 5e-5. During training, we monitored validation accuracy after each epoch, retaining the best-performing model. This process yielded mean validation and test accuracy of 87.5% and 83.6% 8.0%, respectively, and achieved a mean AUROC of 93.1%. For context, human evaluators rated compliance at 86.0% (provided by human evaluators to the original challenges). Thus, our models are close to ground truth accuracy.
For the remaining 2 out of 20 challenges—speaking foreign words (challenge #12) and asking questions (challenge #15)—we assessed compliance through speech transcription (see Table 2).
While we follow machine learning protocols for dataset splitting, we apply the compliance detectors to all identities. Importantly, we do not use any deepfakes for training or validation, as we are only testing compliance.
Content Compliance. We employ Word Information Lost (WIL) (Morris et al., 2004) to quantify the discrepancy between transcribed speech content and the original script. WIL was chosen over standard Word Error Rate (WER) for its emphasis on keywords and de-emphasis of fillers, better reflecting natural conversation dynamics. This metric is applied to all challenges except those involving non-verbal sounds (e.g., coughing or whistling).
@@
Given these three measures, we define a Machine-scored degradation for an audio sample (x) and challenge (chal). A higher score indicates poorer deepfake performance, aligning with our challenge objectives. The degradation score is calculated as:
where indicates compliance with the challenge, returning 1 for compliance and 0 otherwise.
@@
Observations. Regular speech (no-challenge) exhibited low degradation—11% for deepfakes and 5% for originals—indicating RTDFs’ proficiency in replicating standard speech patterns.
@@
Our examination of 20 prospective challenges revealed 10 to be less effective, with AUROC scores below 70% (Table 3), leading to their disqualification. These tasks included variations in accent, reading pace, amplitude, pitch, and specific actions like clapping, coughing/whistling, and nose-holding. Deepfake algorithms demonstrated notable proficiency in mimicking tasks related to regular speech patterns, showcasing the sophistication of voice-cloning systems in replicating human vocal characteristics.
@@
Deepfakes struggled to maintain compliance against top challenges in scenarios requiring lyrical playback, mouth cupping, whispering, and high-pitched voice production (Table 3). These findings indicate a critical vulnerability in current deepfake technologies. Lyrical playback, in particular, introduced measurable vocal distortions, highlighting the effectiveness of these challenges in compromising deepfake audio fidelity.
Our analysis of frequency modulation effects, comparing high-pitch and low-pitch challenges, revealed that high-pitch challenges induced more significant degradation (44%) in deepfake outputs than their low-pitch counterparts (25%). This discrepancy in machine detection efficiency points to limitations in the training datasets of deepfake models, such as the VCTK corpus, which lacks diverse high-pitch voice samples. This limitation underscores the critical role of dataset diversity in enhancing deepfake quality.
Despite the underperformance of specific tasks, we included them in our analysis for comprehensive evaluation and comparison with Yasur et al. (Yasur et al., 2023). The authors had identified six out of the nine ineffective tasks as beneficial against the StarGANv2-VC model (Li et al., 2021). However, these tasks proved less effective when tested against the more sophisticated FREE-VC model (Li et al., 2023b), suggesting that our remaining 10 challenges offer higher degradation even against advanced adversaries.
Takeaways. We have developed a comprehensive and robust machine evaluation framework, identifying 10 tasks that effectively serve as challenges. These challenges enable machines to achieve an AUROC of 88.7%, compared to 56.0% for no-challenge conditions. Our findings highlight the vital role of challenges in identifying deepfake strengths and limitations, particularly when such tasks involve complex auditory challenges beyond simple speech imitation. The analysis reveals two main vulnerabilities of RTDFs: they are restricted by training data diversity, and they do not encode the physiology and physics of human speech apparatus.
Additional Results (Table A.2). We also evaluate two popular deepfake detectors: LFCC feature extraction with LCNN classifier (Liu et al., 2022) and RawNet2 (Tak et al., 2021)/SpecRNet, scoring 75.1% and 86.0% AUC, respectively. Our machine evaluation bypasses them.
Additionally, we evaluate two any-to-many voice converters: StarGANv2-VC (Li et al., 2021) and PPG-VC (Liu et al., 2021). We trained them on our open-sourced dataset to add new identities to their target group. The original implementation of StarGANv2-VC only included 20 speaker identities. In comparison, we trained it on our 100 speaker identities, making them all seen during training. The challenges remain effective as our automated detectors scored AUC of 88.2% and 88.3% against StarGANv2 and PPG, respectively.
@@
Evaluators were tasked with assessing each audio sample’s compliance with the given task () and its overall quality or realism () on a 5-point scale, following the ITU P.808 standard for subjective audio quality assessments (ITU, 2018). See Fig. A.3. †††Reproducible at https://app.gorilla.sc/openmaterials/769577
@@
Here, yields 1 if compliant and 0 otherwise. This measure assigns maximum degradation to a non-compliant sample or a score reduction based on human-rated quality.
Fig. 3 illustrates human evaluation results for the top-10 challenges, while Fig. A.1 illustrates the distribution of human scores across the 20 challenges and the corresponding AUROC.
@@
On the samples corresponding to the top-10 challenges in the balanced dataset, human evaluators achieved a mean AUROC of 87.7% in detecting deepfakes across, comparable to machine evaluations (88.7%), albeit on a difficult, high-quality dataset. However, their accuracy (not AUROC) was lower at 72.6% mean accuracy, in comparison to machines’ 87.7%. See breakdown in first column of Table 5. Interestingly, for this set, the Pearson correlation between human and machine AUROC was merely 0.12, implying a poor positive relationship and revealing the complementary strengths of humans and machines.
@@
| Challenge | Vanilla Acc | Assisted Acc. | Collaborative | ||
@@
✓: correct, ✗: incorrect. Rightmost column: mean SD of confidence change.
@@
Table 5 presents the initial judgments (I), machine-assists (M), and final decisions (F) across the challenges, breaking down the responses across eight possible scenarios and noting changes in confidence levels. When machine predictions aligned with human decisions, they increased human confidence, regardless of correctness. Conversely, when predictions diverged, they reduced human confidence. Notably, when humans initially erred and machines provided correct guidance, humans changed their decisions 45.4% of the time (781 / (781 + 939)). However, when humans were correct and machines erred, the latter misled humans 29.8% of the time (216 / (216 + 510)). This disparity highlights the beneficial role of machines in correcting human errors and their relatively lesser impact in misleading humans when they are correct.
Comparing the presence of rationale against providing no rationale, we observe no significant difference in humans accepting the decision and changing their mistakes (42.4% vs. 43.0%).
Human-AI Collaboration. The machine-assistance framework positioned machines as a warning signal, preventing them from making unilateral final decisions. This approach necessitates convincing humans to correct their initial judgments based on machine input, which could be cumbersome. To bypass this, we introduced a mechanism allowing the machine to override human decisions when it exhibited higher confidence (not correctness of decision).
We computed the machine confidence for each audio sample’s decision using the machine scores from §5.1. The machine confidence was the normalized difference between a predefined threshold () and the sample’s score. Despite machines demonstrating superior accuracy (87.7%), they often exhibited lower confidence than humans. We counteracted this by scaling the machine’s confidence using a temperature parameter.
Fig.5 illustrates the trade-off between retaining human decisions post-machine intervention and the mean accuracy achieved, with Table4 (right part) breaking down the accuracy. At the intersection point in Fig. 5, humans make 56% of decisions (the rest are made by machines), achieving 84.5% accuracy, a performance boost of 16.4% over vanilla evaluation. This balance indicates that a synergistic approach between humans and AI can significantly enhance accuracy while ensuring that humans retain decision-making authority in the majority of cases.
Takeaways. Challenges enable human capabilities to detect deepfake audios, similar to machines, even on a significantly challenging dataset. Comparing human detection to machines reveals similarities and differences depending on the challenge type, suggesting a complementary relationship between human intuition and machine precision. When combined, machine tagging aided in assisting humans where their performance was lacking. Further collaboration between humans and machines resulted in an even greater boost in deepfake detection performance, while keeping humans in control.

### v4

@@ 4.2. Challenges Dataset Curation
We carefully considered the ethical implications of creating and using deepfakes, including potential misuse and privacy concerns. To mitigate these risks, we implemented strict data handling protocols and enforcement to access the generated deepfakes. Our protocols included an Institutional Data Use Agreement restricting usage to approved research purposes, using online storage designed specifically for high-risk data and a participant withdrawal mechanism enabling automated data purging upon request.
@@
We developed the experiments using the Gorilla Experiment Builder (see Fig A.4). The initial data collection involved 100 subjects, equally distributed between smartphone and desktop environments, yielding 18,600 original recordings. This dataset is comparable in scale to the well-known VCTK speech dataset (Yamagishi et al., 2019), which contains 110 speakers with 400 samples each (totaling 44,000 recordings), while offering significantly greater diversity. Our participant pool consisted of individuals aged 18-65, near equal gender distribution, and is multi-racial. This diverse demographic helped ensure a representative sample of speech patterns and characteristics. ***We open source our instrument at https://app.gorilla.sc/openmaterials/722500
@@
Three English-speaking evaluators—a native monolingual, a non-native multilingual, and a tie-breaker—verified each sample. This approach aimed to enhance inclusivity and mitigate demographic bias. We made compliance decisions using majority voting, resulting in an 84% mean compliance across 100 subjects and 89% across challenges.
Choosing RTDF generators. Our analysis revealed that SoTA Text-to-Speech (TTS) models such as TorToise (Betker, 2023) and Bark (suno-ai, 2023) excelled in content modulation but fell short in capturing speech nuances like emotions or prosody, significantly limiting their effectiveness against our challenges. Consequently, we shifted our focus to speech-to-speech systems, which convert source speech into the targeted speaker’s voice while preserving subtle speech elements.
For our evaluation, we selected FREE-VC (Li et al., 2023b) as the primary deepfake generator after assessment of available voice conversion technologies. FREE-VC represents current state-of-the-art in one-shot voice conversion, achieving superior quality (MOS 3.92±0.14 versus StarGANv2-VC’s 3.49±0.11) while requiring only 3-5 seconds of target audio. This aligns with realistic threat scenarios where attackers have limited access to victim voice samples. Its computational efficiency (98ms processing per second of audio) enables real-time conversion necessary for interactive call scenarios, and independent security evaluations (Korshunov et al., 2023) confirm it presents a "high threat" to speaker verification systems with a 78.3% success rate in bypassing authentication (compared to 64.7% for alternatives). As an open-source implementation, FREE-VC represents an accessible tool for potential attackers.
By evaluating our system against this sophisticated model, we ensure challenges effective against FREE-VC will likely work against less advanced voice conversion technologies. However, to prevent relying only on one VC, we had also evaluated StarGANv2-VC (Li et al., 2021) and PPG-VC (Liu et al., 2021) (see Table A.3).
@@ 5. Evaluation
An impersonator’s goal is to mimic a target individual during a call:
By adhering to the challenge in terms of task and content,
@@
Conversely, our challenges aim to disrupt them.
5.1. RQ1. Machine Evaluation
In this subsection, we evaluate the effectiveness of challenges against voice-cloning attacks using an automated authentication system. Our machine evaluation approach is designed to comprehensively assess the effectiveness of challenges against deepfake attempts without requiring reference samples or human interventions. We selected three non-intrusive metrics based on key aspects that determine successful voice impersonation: (1) Challenge Compliance measures the ability to perform the specific requested task, (2) Content Compliance quantifies information preservation in the delivered message, and (3) Realism assesses the naturalness of speech production. These metrics collectively provide a multi-dimensional assessment that captures both technical and perceptual aspects of deepfake quality.
@@
We fine-tuned each Wav2Vec2 model for 30 epochs with a batch size of 64 and a learning rate of 5e-5. During training, we monitored validation accuracy after each epoch, retaining the best-performing model. This process yielded mean validation and test accuracy of 87.5% and 83.6% 8.0%, respectively, and achieved a mean AUROC of 93.1%. For context, human evaluators rated compliance at 86.0% (provided by human evaluators to the original challenges). Thus, our models are close to ground truth accuracy. For details on Wav2Vec2 and further challenge compliance detectors, refer §A.4.
For the remaining 2 out of 20 challenges—speaking foreign words (challenge #12) and asking questions (challenge #15)—we assessed compliance through speech transcription (see Table A.2).
While we follow machine learning protocols for dataset splitting, we apply the compliance detectors to all identities. Importantly, we do not use any deepfakes for training or validation, as we are only testing compliance to the challenge.
Content Compliance. We employ Word Information Lost (WIL) (Morris et al., 2004) to quantify the discrepancy between (a) transcribing the audio sample using WhisperX and (b) the original script. WIL was chosen over standard Word Error Rate (WER) for its emphasis on keywords and de-emphasis of fillers, better reflecting natural conversation dynamics. WIL values range from 0 (perfect preservation) to 1 (complete information loss). This metric is applied to all challenges except those involving non-verbal sounds (e.g., cough/whistle).
@@
Based on these three measures, we define a machine-scored degradation for an audio sample and challenge as follows:
is the binary challenge compliance score (1 if compliant, 0 if non-compliant).
is the Word Information Lost score (higher values indicate more information loss).
is the NISQA realism score on a scale of 1-5. The division by 5 normalizes the realism score to a 0-1 scale.
This balanced formula gives equal weight to all three aspects of deepfake quality assessment. A high degradation score indicates higher degradation (poor performance of deepfake and good performance of challenge) in at least one dimension, which aligns with our objective of identifying challenges that disrupt deepfakes.
These three components represent complementary but distinct dimensions of deepfake performance. Challenge compliance captures whether the deepfake can adapt to specific vocal tasks; content compliance measures whether semantic meaning is preserved; and realism assesses the natural quality of the speech production. While a high-quality deepfake would need to score well in all three dimensions, our findings reveal that different challenges target different vulnerabilities: some primarily affect compliance (e.g., cup mouth), others mainly impact content preservation (e.g., foreign words), and some primarily degrade realism (e.g., whisper).
@@
Results and Observations. Regular speech (no-challenge) exhibited low degradation—11% for deepfakes and 5% for originals—indicating RTDFs’ proficiency in replicating standard speech patterns.
@@
Our examination of 20 prospective challenges revealed 10 to be less effective, with AUROC scores below 70% (grayed in Table 3), leading to their disqualification. These tasks included variations in accent, reading pace, amplitude, pitch, and specific actions like clapping, coughing/whistling, and nose-holding. Deepfake algorithms demonstrated notable proficiency in mimicking tasks related to regular speech patterns, showcasing the sophistication of voice-cloning systems in replicating human vocal characteristics.
@@
Deepfakes struggled to maintain compliance against top challenges in scenarios requiring lyrical playback (#19,#20), mouth cupping (#2), whispering (#3), and high-pitched voice production (#5). Please refer to Table 3 and Fig. 2 (left). These findings indicate a critical vulnerability in current deepfake technologies. Lyrical playback, in particular, introduced measurable vocal distortions, highlighting the effectiveness of these challenges in compromising deepfake audio fidelity.
Our analysis of frequency modulation effects, comparing high-pitch and low-pitch challenges, revealed that high-pitch (#5) challenges induced more significant degradation (44%) in deepfake outputs than their low-pitch (#6) counterparts (25%). This discrepancy in machine detection efficiency points to limitations in the training datasets of deepfake models, such as the VCTK corpus, which lacks diverse high-pitch voice samples. This limitation underscores the critical role of dataset diversity in enhancing deepfake quality.
In addition to FREE-VC (Li et al., 2023b), we also evaluated two any-to-many voice converters: StarGANv2-VC (Li et al., 2021) and PPG-VC (Liu et al., 2021). We trained them on our open-sourced dataset to add new identities to their target group. The original implementation of StarGANv2-VC only included 20 speaker identities. In comparison, we trained it on our 100 speaker identities, making them all seen during training. The challenges remain effective as our automated detectors scored AUC of 88.2% and 88.3% against StarGANv2 and PPG, respectively.
In comparison to prior work, D-CAPTCHA (Yasur et al., 2023), our work identified that six out of their nine tasks were ineffective. As they had evaluated only against StarGANv2-VC (Li et al., 2021), these tasks proved less effective when tested against the more sophisticated FREE-VC model (Li et al., 2023b), suggesting that our remaining 10 challenges offer higher degradation even against advanced adversaries.
Key Takeaway. We have developed a comprehensive and robust machine evaluation framework, identifying 10 tasks that effectively serve as challenges. These challenges enable machines to achieve an AUROC of 88.7%, compared to 56.0% for no-challenge conditions. Our findings highlight the vital role of challenges in identifying deepfake strengths and limitations, particularly when such tasks involve complex auditory challenges beyond simple speech imitation. The analysis reveals two main vulnerabilities of RTDFs: they are restricted by training data diversity, and they do not encode the physiology and physics of human speech apparatus.
@@
Evaluators were tasked with assessing each audio sample’s compliance with the given task () and its overall quality or realism () on a 5-point Likert scale, following the ITU P.808 standard for subjective audio quality assessments (ITU, 2018) (see Fig. A.3).
@@
Here, is 1 if human response was compliant and 0 otherwise. Also is a realism rating on a 5-point scale (higher is better). This measure assigns maximum degradation to a non-compliant sample or a score reduction based on human ratings.
Fig. 2 (right) illustrates human evaluation results for the top-10 challenges, while Fig. A.2 illustrates the distribution of human scores across the 20 challenges and the corresponding AUROC.
@@
On the samples corresponding to the top-10 machine-validated challenges in the balanced dataset, human evaluators achieved a mean AUROC of 87.7% in detecting deepfakes across, comparable to machine evaluations (88.7%), albeit on a difficult, high-quality dataset. However, their accuracy (not AUROC) was lower at 72.6% mean accuracy, in comparison to machines’ 87.7%. See breakdown in Table 4 (left). Interestingly, for this set, the Pearson correlation between human and machine AUROC was merely 0.12, implying a poor positive relationship and revealing the complementary strengths of humans and machines.
@@
| Challenge | Vanilla Acc. | Assisted Acc. | Collaborative | ||
@@
Table 5 presents the initial judgments (I), machine-assists (M), and final decisions (F) across the challenges, breaking down the responses across eight possible scenarios and noting changes in confidence levels. When machine predictions aligned with human decisions, they increased human confidence, regardless of correctness. Conversely, when predictions diverged, they reduced human confidence. Notably, when humans initially erred and machines provided correct guidance, humans changed their decisions 45.4% of the time (781 / (781 + 939)). However, when humans were correct and machines erred, the latter misled humans 29.8% of the time (216 / (216 + 510)). This disparity highlights the beneficial role of machines in correcting human errors and their relatively lesser impact in misleading humans when they are correct. We observe no significant difference (42.4% vs. 43.0%) in humans accepting the decision when given the rationale or otherwise.
Human-AI Collaboration. Our framework implements a confidence based decision routing system that determines when human review is necessary or when machine decision is acceptable. We used a detection threshold for classifying samples as deepfakes, accepted across literature (Ravanelli et al., 2021). For each audio sample, machine confidence is computed as , representing the normalized distance from the decision boundary.
To address the discrepancy between high machine accuracy (87.7%) and its low confidence, we introduced a temperature parameter that calibrates confidence scaling: . This calibration determines which decisions can be automated versus requiring human review. Calibrated samples with high confidence () receive automated decisions, while low-confidence cases () are routed to human reviewers with machine predictions, such as ‘Deepfake-Likely,’ displayed as assistive information.
Fig. 4 illustrates this accuracy-automation trade-off across temperature values. At the optimal operating point (), humans make 56% of decisions while machines handle the remaining 44%, achieving 84.5% accuracy—a 16.4% improvement over human-only evaluation (refer Table 4 (right)). This balanced approach ensures human oversight for uncertain cases while leveraging machine precision for high-confidence decisions.
@@
Note: I: Initial decision, M: Machine verdict, F: Final decision. ✓: correct, ✗: incorrect.
Key Takeaway. Challenges enable human capabilities to detect deepfake audios, similar to machines, even on a significantly challenging dataset. Comparing human detection to machines reveals similarities and differences depending on the challenge type, suggesting a complementary relationship between human intuition and machine precision. When combined, machine tagging aided in assisting humans where their performance was lacking. Further collaboration between humans and machines resulted in an even greater boost in deepfake detection performance, while keeping humans in control.